what is the first five terms and 50th term of this sequence.
An=2a n-1 + 5 and a1=3

Found 3 solutions by ikleyn, greenestamps, Edwin McCravy:
Answer by ikleyn(52776)

.


In this problem, you are given a recurrent formula to calculate next term of the sequence

using its previous step.  You also are given the first term.


So, to get the second term, simply substitute first term a1 = 3 into the formula

    a2 = 2*3 + 5 = 6 + 5 = 11.


To get the third term, substitute the second term a2 = 11 into the formula

    a3 = 2*11 + 5 =22 + 5 = 27.


Continue doing this way until you will get a5.



Answer by greenestamps(13198)



The recursive formula says we get the next term in the sequence by doubling the current term and adding 5. Starting with 3 as the first term...

a(1) = 3
a(2) = 3(2)+5 = 11
a(3) = 11(2)+5 = 27
a(4) = 27(2)+5 = 59
a(5) = 59(2)+5 = 123

It would be very tedious to find the 50th term by continuing to use the recursive formula, so to find the 50th term we would like to find an explicit formula for the n-th term.

To do that, we could try the method of finite differences. Here are the first few terms of the sequence and the first and second differences:

    3   11    27   59  123
      8    16   32   64
         8    16   32

That pattern of differences tells us that the formula for the n-th term is not a polynomial but instead is based on powers of 2.

Playing with the first few numbers in the sequence, we see that adding 5 to each term gives us a sequence of powers of 2:

8, 16, 32, 64, 128, ...

So...
a(1) is 8-5 = 2^3-5
a(2) is 16-5 = 2^4-5
a(3) is 32-5 = 2^5-5
...

And we see the pattern for the n-th term is

a(n) = 2^(n+2)-5

So the 50th term is

2^52-5 = 9007199254740987

My TI-84 calculator won't calculate that number; and I tried using excel but it fell one digit short of being able to calculate the exact number. I used an online program (pari) to find the exact answer.


Answer by Edwin McCravy(20054)

I wrote a program in LibertyBasic using the given recursion formula
and greenestamps' general formula.  Here is the LibertyBasic progrem:

  for n=1 to 50
  if n=1 then a=3: goto 1
  a=2*a+5
1 print n, a, 2^(n+2)-5
  next

They are identical, as you can see from the complete output below.  However, the
50th term does not agree with greenestamps' 50th term.  He made a slight
calculator error and got the 51st term by calculating 2^53-5 instead of 2^52-5.

              2a(n-1)-5
              a(1)=3                      a(n)=2n+2-5                                                   
1             3                           3
2             11                          11
3             27                          27
4             59                          59
5             123                         123
6             251                         251
7             507                         507
8             1019                        1019
9             2043                        2043
10            4091                        4091
11            8187                        8187
12            16379                       16379
13            32763                       32763
14            65531                       65531
15            131067                      131067
16            262139                      262139
17            524283                      524283
18            1048571                     1048571
19            2097147                     2097147
20            4194299                     4194299
21            8388603                     8388603
22            16777211                    16777211
23            33554427                    33554427
24            67108859                    67108859
25            134217723                   134217723
26            268435451                   268435451
27            536870907                   536870907
28            1073741819                  1073741819
29            2147483643                  2147483643
30            4294967291                  4294967291
31            8589934587                  8589934587
32            17179869179                 17179869179
33            34359738363                 34359738363
34            68719476731                 68719476731
35            137438953467                137438953467
36            274877906939                274877906939
37            549755813883                549755813883
38            1099511627771               1099511627771
39            2199023255547               2199023255547
40            4398046511099               4398046511099
41            8796093022203               8796093022203
42            17592186044411              17592186044411
43            35184372088827              35184372088827
44            70368744177659              70368744177659
45            140737488355323             140737488355323
46            281474976710651             281474976710651
47            562949953421307             562949953421307
48            1125899906842619            1125899906842619
49            2251799813685243            2251799813685243
50            4503599627370491            4503599627370491

Edwin